# Deep Transformer Network for Monocular Pose Estimation of Ship-Based UAV

### v1

@@ Abstract
This paper introduces a deep transformer network for estimating the relative 6D pose of a Unmanned Aerial Vehicle (UAV) with respect to a ship using monocular images. A synthetic dataset of ship images is created and annotated with 2D keypoints of multiple ship parts. A Transformer Neural Network model is trained to detect these keypoints and estimate the 6D pose of each part. The estimates are integrated using Bayesian fusion. The model is tested on synthetic data and in-situ flight experiments, demonstrating robustness and accuracy in various lighting conditions. The position estimation error is approximately 0.8% and 1.0% of the distance to the ship for the synthetic data and the flight experiments, respectively. The method has potential applications for ship-based autonomous UAV landing and navigation. Source code : fdcl-gwu/TNN-MO, Video : https://youtu.be/ZG_zVVS8xw8
@@ 1 Introduction
Unmanned Aerial Vehicles (UAVs) have seen a surge in usage across a multitude of industries, such as aerial photography, military operations, agriculture, mapping, and surveying. The advantages of UAVs over traditional manned aircraft are numerous, including cost-effectiveness, enhanced safety, and superior flexibility. However, the autonomous operation of UAVs, particularly their ability to land on moving platforms like ships, poses a crucial challenge. This capability is of significant importance for industries that depend on maritime transportation or offshore operations.
A primary challenge in this context is the estimation of the UAV’s relative pose with respect to the ship, which is vital for precise control of the UAV’s movements and ensuring a safe landing. Conventionally, the relative pose has been determined using the Real-Time Kinematic (RTK) Global Positioning System (GPS). To receive RTK-GPS, a communication link between the ship and the UAV must be maintained at all times, typically via radio. As a result, the UAV cannot estimate its relative position independently. In military surveillance operations, the broadcasting of radio signals to maintain this communication link could potentially reveal the ship’s position, thereby increasing its vulnerability. Also, reliance on GPS can lead to issues in situations where GPS signals may be weak or unavailable, and it is susceptible to malicious activities, such as jamming or spoofing, which can disrupt the operation of the UAV.
An alternative to GPS for estimating the relative pose of a UAV could involve the use of a monocular camera mounted on the UAV. Then, captured Red-Green-Blue (RGB) camera images can be processed to estimate the UAV’s relative pose. This method provides a potential solution for independent and secure pose estimation. Traditionally, ArUco markers have been employed to estimate the relative pose of the camera. ArUco markers are a type of fiducial marker, which are square markers with a specific binary pattern that can be easily detected and identified by a computer [1, 2]. However, the use of ArUco markers comes with certain disadvantages, such as the need for clear visibility and specific lighting conditions, their susceptibility to occlusion, and the requirement for the markers to be within the camera’s field of view. These limitations can affect the accuracy and reliability of pose estimation.
@@
To address this issue, in this paper, we propose an innovative method that facilitates multi-object detection using the Transformer Neural Network (TNN). In this approach, the ship is decomposed into multiple parts with varying sizes and locations, as opposed to the previous single object detection method. We model a Transformer Neural Network for Multi Object (TNN-MO) by utilizing the DETR architecture to estimate the 6D pose of a UAV relative to the selected parts of the ship. The multiple pose estimates with respect to those parts are integrated in a Bayesian fashion.
@@ 2.1 Development of the 3D CAD Model
First, we create a CAD model by capturing point clouds of the vessel. We captured a footage of the entire ship using a monocular RGB camera mounted on an octocopter [10] [11]. Utilizing this video, we generated a 3D point cloud of the ship using structure from motion techniques in computer vision [12]. The resulting 3D point cloud is corrupted by noise and it incomplete as the octocopter is not allowed to fly near the bow side of the ship. It is first was manually filtered to remove outliers, then the cleaned point cloud was imported into an open-source rendering software, namely Blender, to create a CAD model [13]. During this step, detailed structures on the ship not captured by the point cloud, such as the ladders, ship bow, and fence, were manually added too. Furthermore, the CAD model was rescaled according to the actual size of the ship. To enhance the realistic view of the scene, we included other objects such as the ocean, sun, and humans. Also, to make ocean more realistic, we randomly generated wave patterns for every scene using Blender wave modifier tool.
Figure 2: 3D model based on the filtered point cloud and a CAD model obtained from the point cloud
@@ 2.2 Assignment of Textures
To address this issue, we generate synthetic images using domain randomization, where non-photo-realistic and realistic textures are randomly assigned to each object in Blender [14]. Texture assignment, which involves mapping necessary textures onto 3D objects, is a crucial step in the process of generating synthetic images. In this work, we utilized two sources to create texture categories: the Describable Textures Dataset (DTD) [15] and the open-source platform OpenGameArt [16]. The categorization of these textures was based on their image frequency and the characteristics of the textures. Specifically, textures with high frequencies or those resembling oceans were grouped together as "Ocean-like textures". Textures selected with low frequencies were categorized as "Vessel skin textures". Similarly, textures with low frequencies or those resembling the sky were grouped together as "Sky-like textures". To define the landing pad and its markings, we grouped together dark, asphalt-like textures as "Landing pad textures", while whitish color textures were categorized as "Landing pad markings textures", as shown in Figure 3.
@@ 2.3 Environmental Illumination and Occlusions
Figure 4: Illustration of randomly varied ambient lighting conditions and human models
In the ocean environments, the lighting conditions may vary drastically. Furthermore, the illumination of objects in an image can be affected by changes in the orientation of the objects or the camera, even under consistent lighting conditions. For instance, the orientation of an object or the camera may result in the appearance of shadows under direct sunlight. Most cameras are equipped with an automatic exposure control (AE) system that adjusts the brightness of the image based on the lighting conditions. As a result, the lighting conditions in the captured image can vary depending on the adjustments made by the camera’s built-in AE system. This variability in lighting conditions poses challenges for object detection, tracking, and pose estimation when using real images. As such, randomizing object texture may not fully mitigate these challenges. To address this, we have enhanced the scene by introducing randomly varying lighting conditions, as illustrated in Figure 4.
@@ 3.1 Formulating Probability Density for Position
We first present the probability distribution for $C$ and $F$. Let the base frame $B$ be fixed to the flight deck of the ship. The $x$-axis of this frame points to the starboard side of the ship (right side of the helmsman), the $y$-axis points toward the bow (front) of the ship, and the $z$-axis points upwards. These axes are denoted by $x_{B}$, $y_{B}$, and $z_{B}$, as illustrated in Figure 5. The location of $C$ or $F$ is defined by the spherical coordinates $(r,\theta_{B},\phi_{B})$, where $r\in\mathbb{R}$ represents the distance, and $\theta_{B}\in[-\pi,\pi],\phi_{B}\in[0,\frac{\pi}{2}]$ represent rotation around the $z$-axis and rotation around the $x$-axis respectively.
For the position of the camera $C$, the angle $\theta_{B}$ is sampled from a uniform distribution on the range $[-\pi,0]$, representing the area behind the flight deck. Similarly, $\phi_{B}$ is uniformly sampled from the range $[0,\frac{\pi}{3}]$, corresponding to the area above the flight deck. This choice is motivated by our desire to generate realistic camera orientations that resemble those of a horizontally fixed camera during UAV flight. We initially experimented with increasing the maximum range of $\phi_{B}$ closer to $\frac{\pi}{2}$, but found that this led to a number of unrealistic camera orientations. To mitigate this, we decided to set the maximum range for $\phi_{B}$ at $\frac{\pi}{3}$. This decision has proven effective in maintaining the realism of our generated camera orientations. The distance $r$ is sampled from a truncated normal distribution, which is a one-dimensional Gaussian distribution truncated to the range of $[0,L]$ and rescaled accordingly. Here, the maximum range $L$ is chosen as $25m$, and the mean and the standard deviation of the distribution are $1$ and $40$, respectively. Next, the distribution for the target point $F$ is sampled in a similar manner. The range for each of $\theta_{B}$, $\phi_{B}$, and $r$ is $[0,2\pi]$, $[-\frac{\pi}{2},\frac{\pi}{2}]$, and $[0,15]$, respectively. The mean and the standard deviation for the truncated normal distribution are 00 and $1$. The resulting set of sampled $C$ and $F$ is presented at Figure 6.
@@
Figure 5: The base frame and the camera frame
@@ 3.2 Attitude of Camera
Once the points $C$ and $F$ are sampled, the camera attitude is determined as follows. Let the camera frame be defined such that its origin coincides with $C$, and the positive x-axis points to the right (along the direction of the image width), the positive y-axis points upward (along the direction of the image height), and the positive z-axis is opposite to the line of sight, as illustrated in Figure 5. The orientation of the camera frame with respect to the flight deck frame is specified by a rotation matrix $R\in\mathsf{SO(3)}=\{R\in\mathbb{R}^{3\times 3}\,|\,R^{T}R=I,\;\mathrm{det}[R]=1\}$.
@@ 3.3 Generated Synthetic Images
The selected camera pose is transferred to the rendering software, Blender to generate the synthetic images with varying textures, wave patterns, and lightening conditions. Each image is paired with the correct pose $(C,R)$ to be used for training and verification. The selected samples from the synthetic images are illustrated in Figure 7.
Figure 7: Generated synthetic images with the presence of sea horizon
@@
To address this, we have opted to generate a distinct synthetic image dataset without the horizon. This is achieved by substituting the ship’s background with a randomly generated image. The random image is produced using the random image generator as proposed by [17]. Then, we mix the above dataset with the synthetic images without any horizon. This approach ensures a more diverse and realistic representation in our synthetic image dataset, while capturing the importation information provided by a horizon.
Figure 8: Generated synthetic images without sea horizon by replacing background with random image
@@ 3.4 Decomposition of the Ship into Multiple Parts and Keypoint Selection
However, defining the entire ship as a single object is not desirable, as the visibility might be obscured depending on the position and the orientation of the camera. For example, during the landing, the keypoints at the stern of the ship may not be visible. To address this issue, the ship is decomposed into multiple parts for detection, rather than being detected as a whole. This method also helps mitigate the impact of certain camera orientations and specific lighting conditions that might result in only partial visibility of certain areas of the ship, potentially affecting the accuracy of the model’s estimations. Specifically, the ship is decomposed into six parts: the whole ship, stern, superstructure, and three parts of the flight structure immediately forward of the flight deck, referred to as the dog house,, as illustrated at Figure 9.
Figure 9: Selected specific components of the ship and 3D keypoints
@@ 4.1 Transformer Neural Network for Keypoints Estimation
Refer to captionRefer to captionCNNEPnPEPnPEPnPEPnPBackbonePositional EncodingObject QueriesPrediction HeadsFeatures transformer
@@
Figure 10: TNN-MO model architecture to estimate 6D pose
The proposed transformer neural network architecture is described as follows. First, for a given RGB image, we extract image features using a CNN backbone. Specifically, the input RGB image has a size of $H\times W=480\times 640$. We employ ResNet50 to generate a lower-resolution feature $f\in\mathbb{R}^{C\times H_{0}\times W_{0}}$, where $C=2048$ is the channel dimension and $H_{0},W_{0}$ are defined as $\frac{H}{32}=15,\frac{W}{32}=20$ respectively, following the DETR [6]. Then, we use a $1\times 1$ convolution to reduce the feature dimension to $d=256$. The extracted features, which are in the dimension of $d\times H_{0}\times W_{0}$, are flattened to a dimension of $d\times H_{0}W_{0}$. Subsequently, a fixed positional encoding is added to the features to generate input embeddings, which have the dimension of $d\times H_{0}W_{0}$. Then, each of these embeddings is used as an input for the transformer encoder [18].
The transformer encoder has six standard encoder layers with skip connections. The output from the encoder is passed to the decoder module which also has six standard decoder layers with skip connections, accompanied by $Q$ object queries [7]. These object queries are a set of learnable positional embeddings with the dimension $d\times Q$, representing the position of each object class. These vectors are learned during the training process of the model, to help the model understand and represent the positional relationships between different objects in the data. Let $N$ be the cardinality of the object classes including the case when no object is detected, which is denoted by $\varnothing$. Since we are considering six parts of the ship as objects which defined in Figure 9, we have $N=7$. In the presented TNN-MO model, we set $Q=N$, implying that the number of object queries is equal to the object classes.
@@ 4.2 Ground Truth Data
In the training input image, the ground truth object class labels are represented by $\mathbf{c}_{g}=[c_{g_{1}},c_{g_{2}},\ldots,c_{g_{N}}]\in\{0,1\}^{N}$, where $c_{g_{j}}$ is a binary variable indicating the presence (or the absence) of the $j$-th object. Specifically, if the $j$-th object appears in the training input image, we set $c_{g_{j}}=1$; otherwise, we set $c_{g_{j}}=0$. The $N$-th object corresponds to the $\varnothing$ object, and therefore, $c_{g_{N}}=0$ always for any image in the training data set. It is important to note that since we are considering a single ship, there is no repetition of objects in the image. For example, following the object classes defined in Figure 9, when the whole ship (5th) and the ship stern (6th) are visible in a specific image, the resulting ground truth object class labels is $\mathbf{c}_{g}=[0,0,0,0,1,1,0]$.

### v2

# Deep Transformer Network for Monocular Pose Estimation of Shipborne Unmanned Aerial Vehicle

@@ Abstract
This paper introduces a deep transformer network for estimating the relative 6D pose of a Unmanned Aerial Vehicle (UAV) with respect to a ship using monocular images. A synthetic dataset of ship images is created and annotated with 2D keypoints of multiple ship parts. A Transformer Neural Network model is trained to detect these keypoints and estimate the 6D pose of each part. The estimates are integrated using Bayesian fusion. The model is tested on synthetic data and in-situ flight experiments, demonstrating robustness and accuracy in various lighting conditions. The position estimation error is approximately 0.8% and 1.0% of the distance to the ship for the synthetic data and the flight experiments, respectively. The method has potential applications for ship-based autonomous UAV landing and navigation.
33footnotetext: Preliminary results of this paper were presented at the 2024 AIAA SciTech Forum, held at the Hyatt Regency Orlando, 9801 International Drive, Orlando, Florida, USA, from January 8 to January 12, 2024; AIAA Paper Number: AIAA-2024-2877.
@@ 1 Introduction
Unmanned Aerial Vehicles (UAVs) have seen a surge in usage across a multitude of industries, such as aerial photography, military operations, agriculture, mapping, and surveying. The advantages of UAVs over traditional manned aircraft are numerous, including cost-effectiveness, enhanced safety, and superior flexibility. However, the autonomous operation of UAVs, particularly their ability to land on moving platforms like ships, poses a crucial challenge [1]. The unique characteristics of maritime environments—constant motion, unpredictable wave patterns, and rapidly changing atmospheric and light conditions—significantly amplify the complexity of autonomous landing. This challenge is particularly critical in maritime and naval operations, where precision and adaptability are paramount. This capability is of significant importance for industries that depend on maritime transportation or offshore operations.
A primary challenge in this context is the estimation of the UAV’s relative pose with respect to the ship, which is vital for precise control of the UAV’s movements and ensuring a safe landing. Conventionally, the relative pose has been determined using the Real-Time Kinematic (RTK) Global Positioning System (GPS). The RTK-GPS provides highly accurate positioning by using differential corrections to account for errors in GPS signals. To receive RTK-GPS, a communication link between the ship and the UAV must be maintained at all times, typically via radio [2]. Consequently, the UAV relies on the additional measurements on the ship, which limit its ability to independently determine its relative position. In military surveillance operations, the broadcasting of radio signals to maintain this communication link could potentially reveal the ship’s position, thereby increasing its vulnerability. Also, reliance on GPS can lead to issues in situations where GPS signals may be weak or unavailable, and it is susceptible to malicious activities, such as jamming or spoofing, which can disrupt the operation of the UAV.
An alternative to GPS for estimating the relative pose of a UAV could involve the use of a monocular camera mounted on the UAV, where the captured images can be processed to estimate the UAV’s relative pose. Detecting fiducial markers such as lights, laser pointers, or specific visual patterns using cameras are well understood in computer vision and they are often used for position and attitude estimation during landing operations [3, 4]. For example, ArUco markers are composed of square patterns with specific binary codes that can be easily detected and identified by a computer [5].
They have been used to estimate the UAV’s relative pose with respect to the ship [6, 7]. However, the use of fiducial markers comes with certain disadvantages, such as the need for clear visibility and specific lighting conditions, their susceptibility to occlusion, and the requirement for the markers to be within the camera’s field of view. These limitations can affect the accuracy and reliability of pose estimation.
@@
To address this issue, in this paper, we propose an innovative method that facilitates multi-object detection using the Transformer Neural Network (TNN)  [^4] Source code : fdcl-gwu/TNN-MO, Video : https://youtu.be/ZG_zVVS8xw8 as shown in Fig. 1. In this approach, the ship is decomposed into multiple parts with varying sizes and locations, as opposed to the previous single object detection method. We model a Transformer Neural Network for Multi Object (TNN-MO) by utilizing the DETR architecture to estimate the 6D pose of a UAV relative to the selected parts of the ship. The multiple pose estimates with respect to those parts are integrated in a Bayesian fashion.
@@ 2.1 Development of the 3D CAD Model
First, we create a CAD model of the ship by capturing point clouds of the vessel as shown Fig. 2. We captured a footage of the entire ship using a monocular RGB camera mounted on an octocopter [15] [16]. Utilizing this video, we generated a 3D point cloud of the ship using structure from motion techniques in computer vision [17]. The resulting 3D point cloud is corrupted by noise and it incomplete as the octocopter is not allowed to fly near the bow side of the ship. It is first was manually filtered to remove outliers, then the cleaned point cloud was imported into an open-source rendering software, namely Blender, to create a CAD model [18]. During this step, detailed structures on the ship not captured by the point cloud, such as the ladders, ship bow, and fence, were manually added too. Furthermore, the CAD model was rescaled according to the actual size of the ship. To enhance the realistic view of the scene, we included other objects such as the ocean, sun, and humans. Also, to make ocean more realistic, we randomly generated wave patterns for every scene using Blender wave modifier tool.
Figure 2: 3D model based on the filtered point cloud and a CAD model obtained from the point cloud.
@@ 2.2 Assignment of Textures
To address this issue, we generate synthetic images using domain randomization, where non-photo-realistic and realistic textures are randomly assigned to each object in Blender [19]. Texture assignment, which involves mapping necessary textures onto 3D objects, is a crucial step in the process of generating synthetic images. In this work, we utilized two sources to create texture categories: the Describable Textures Dataset (DTD) [20] and the open-source platform OpenGameArt [21]. The categorization of these textures was based on their image frequency and the characteristics of the textures. Specifically, textures with high frequencies or those resembling oceans were grouped together as "Ocean-like textures". Textures selected with low frequencies were categorized as "Vessel skin textures". Similarly, textures with low frequencies or those resembling the sky were grouped together as "Sky-like textures". To define the landing pad and its markings, we grouped together dark, asphalt-like textures as "Landing pad textures", while whitish color textures were categorized as "Landing pad markings textures", as shown in Fig. 3.
@@ 2.3 Environmental Illumination and Occlusions
In the ocean environments, the lighting conditions may vary drastically. Furthermore, the illumination of objects in an image can be affected by changes in the orientation of the objects or the camera, even under consistent lighting conditions. For instance, the orientation of an object or the camera may result in the appearance of shadows under direct sunlight. Most cameras are equipped with an automatic exposure control (AE) system that adjusts the brightness of the image based on the lighting conditions. As a result, the lighting conditions in the captured image can vary depending on the adjustments made by the camera’s built-in AE system. This variability in lighting conditions poses challenges for object detection, tracking, and pose estimation when using real images. As such, randomizing object texture may not fully mitigate these challenges. To address this, we have enhanced the scene by introducing randomly varying lighting conditions, as illustrated in Fig. 4.
Figure 4: Illustration of randomly varied ambient lighting conditions and human models.
@@ 3.1 Formulating Probability Density for Position
We first present the probability distribution for $C$ and $F$. Let the base frame $B$ be fixed to the flight deck of the ship. The $x$-axis of this frame points to the starboard side of the ship (right side of the helmsman), the $y$-axis points toward the bow (front) of the ship, and the $z$-axis points upwards. These axes are denoted by $x_{B}$, $y_{B}$, and $z_{B}$, as illustrated in Fig. 5. The location of $C$ or $F$ is defined by the spherical coordinates $(r,\theta_{B},\phi_{B})$, where $r\in\mathbb{R}$ represents the distance, and $\theta_{B}\in[-\pi,\pi],\phi_{B}\in[0,\frac{\pi}{2}]$ represent rotation around the $z$-axis and rotation around the $x$-axis respectively.
For the position of the camera $C$, the angle $\theta_{B}$ is sampled from a uniform distribution on the range $[-\pi,0]$, representing the area behind the flight deck. Similarly, $\phi_{B}$ is uniformly sampled from the range $[0,\frac{\pi}{3}]$, corresponding to the area above the flight deck. This choice is motivated by our desire to generate realistic camera orientations that resemble those of a horizontally fixed camera during UAV flight. We initially experimented with increasing the maximum range of $\phi_{B}$ closer to $\frac{\pi}{2}$, but found that this led to a number of unrealistic camera orientations. To mitigate this, we decided to set the maximum range for $\phi_{B}$ at $\frac{\pi}{3}$. This decision has proven effective in maintaining the realism of our generated camera orientations. The distance $r$ is sampled from a truncated normal distribution, which is a one-dimensional Gaussian distribution truncated to the range of $[0,L]$ and rescaled accordingly. Here, the maximum range $L$ is chosen as $25m$, and the mean and the standard deviation of the distribution are $1$ and $40$, respectively. Next, the distribution for the target point $F$ is sampled in a similar manner. The range for each of $\theta_{B}$, $\phi_{B}$, and $r$ is $[0,2\pi]$, $[-\frac{\pi}{2},\frac{\pi}{2}]$, and $[0,15]$, respectively. The mean and the standard deviation for the truncated normal distribution are 00 and $1$. The resulting set of sampled $C$ and $F$ is presented at Fig. 6.
@@
Figure 5: The base frame and the camera frame.
@@ 3.2 Attitude of Camera
Once the points $C$ and $F$ are sampled, the camera attitude is determined as follows. Let the camera frame be defined such that its origin coincides with $C$, and the positive x-axis points to the right (along the direction of the image width), the positive y-axis points upward (along the direction of the image height), and the positive z-axis is opposite to the line of sight, as illustrated in Fig. 5. The orientation of the camera frame with respect to the flight deck frame is specified by a rotation matrix $R\in\mathsf{SO(3)}=\{R\in\mathbb{R}^{3\times 3}\,|\,R^{T}R=I,\;\mathrm{det}[R]=1\}$.
@@ 3.3 Generated Synthetic Images
The selected camera pose is transferred to the rendering software, Blender to generate the synthetic images with varying textures, wave patterns, and lightening conditions. Each image is paired with the correct pose $(C,R)$ to be used for training and verification. The selected samples from the synthetic images are illustrated in Fig. 7.
Figure 7: Generated synthetic images with the presence of sea horizon.
@@
To address this, we have opted to generate a distinct synthetic image dataset without the horizon. This is achieved by substituting the ship’s background with a randomly generated image. The random image is produced using the random image generator as proposed by [22]. Then, we mix the above dataset with the synthetic images without any horizon as shown in Fig. 8. This approach ensures a more diverse and realistic representation in our synthetic image dataset, while capturing the importation information provided by a horizon.
Figure 8: Generated synthetic images without sea horizon by replacing background with random image.
@@ 3.4 Decomposition of the Ship into Multiple Parts and Keypoint Selection
However, defining the entire ship as a single object is not desirable, as the visibility might be obscured depending on the position and the orientation of the camera. For example, during the landing, the keypoints at the stern of the ship may not be visible. To address this issue, the ship is decomposed into multiple parts for detection, rather than being detected as a whole. This method also helps mitigate the impact of certain camera orientations and specific lighting conditions that might result in only partial visibility of certain areas of the ship, potentially affecting the accuracy of the model’s estimations. Specifically, the ship is decomposed into six parts: the whole ship, stern, superstructure, and three parts of the flight structure immediately forward of the flight deck, referred to as the dog house,, as illustrated at Fig. 9.
Figure 9: Selected specific components of the ship and 3D keypoints.
@@ 4.1 Transformer Neural Network for Keypoints Estimation
Refer to captionPositional EncodingRefer to captionCNNEPnPEPnPEPnPEPnPBackboneObject QueriesPrediction HeadsFeatures transformer
@@
Figure 10: TNN-MO model architecture to estimate 6D pose.
The proposed transformer neural network architecture, as shown in Fig. 10, is described as follows. First, for a given RGB image, we extract image features using a CNN backbone. Specifically, the input RGB image has a size of $H\times W=480\times 640$. We employ ResNet50 to generate a lower-resolution feature $f\in\mathbb{R}^{C\times H_{0}\times W_{0}}$, where $C=2048$ is the channel dimension and $H_{0},W_{0}$ are defined as $\frac{H}{32}=15,\frac{W}{32}=20$ respectively, following the DETR [11]. Then, we use a $1\times 1$ convolution to reduce the feature dimension to $d=256$. The extracted features, which are in the dimension of $d\times H_{0}\times W_{0}$, are flattened to a dimension of $d\times H_{0}W_{0}$. Subsequently, a fixed positional encoding is added to the features to generate input embeddings, which have the dimension of $d\times H_{0}W_{0}$. Then, each of these embeddings is used as an input for the transformer encoder [23].
The transformer encoder has six standard encoder layers with skip connections. The output from the encoder is passed to the decoder module which also has six standard decoder layers with skip connections, accompanied by $Q$ object queries [12]. These object queries are a set of learnable positional embeddings with the dimension $d\times Q$, representing the position of each object class. These vectors are learned during the training process of the model, to help the model understand and represent the positional relationships between different objects in the data. Let $N$ be the cardinality of the object classes including the case when no object is detected, which is denoted by $\varnothing$. Since we are considering six parts of the ship as objects which defined in Fig. 9, we have $N=7$. In the presented TNN-MO model, we set $Q=N$, implying that the number of object queries is equal to the object classes.
@@ 4.2 Ground Truth Data
In the training input image, the ground truth object class labels are represented by $\mathbf{c}_{g}=[c_{g_{1}},c_{g_{2}},\ldots,c_{g_{N}}]\in\{0,1\}^{N}$, where $c_{g_{j}}$ is a binary variable indicating the presence (or the absence) of the $j$-th object. Specifically, if the $j$-th object appears in the training input image, we set $c_{g_{j}}=1$; otherwise, we set $c_{g_{j}}=0$. The $N$-th object corresponds to the $\varnothing$ object, and therefore, $c_{g_{N}}=0$ always for any image in the training data set. It is important to note that since we are considering a single ship, there is no repetition of objects in the image. For example, following the object classes defined in Fig. 9, when the whole ship (5th) and the ship stern (6th) are visible in a specific image, the resulting ground truth object class labels is $\mathbf{c}_{g}=[0,0,0,0,1,1,0]$.